f(x)= (2x -1) / (x-3)
Find the inverse function.
I get to y(x-3)=2x-1 and from there the algebra confuses me. I can't get one variable isolated.

Answer by stanbon(75887)

f(x)= (2x -1) / (x-3)
Find the inverse function.
-----
1st: Interchange x and y to get::
x = (2y-1)/(y-3)
-----
2nd: Solve for "y":
xy-3x = 2y-1
-----
2y-xy = 1-3x
------
y(2-x) = 1-3x
------
y = (3x-1)/(x-2)
======================
Cheers,
Stan H.
======================